Solve the differential equation by substitution suggested by equation. Show complete solution.





(5x+3e^y)dx+2xe^y dy =0

1
Expert's answer
2021-11-07T17:34:27-0500

"(5x+3e^y)dx+2xe^y dy =0\\
2xe^y dy=-((5x+3e^y)dx)\\
e^y \frac{dy}{dx}=-(\frac{5x+3e^y}{2x})"

Let us substitute "e^y=z\Rightarrow e^y \frac{dy}{dx}=\frac{dz}{dx}\\"

Then, "\frac{dz}{dx}=-(\frac{5x+3e^y}{2x})"

"\Rightarrow \frac{d z}{d x}=-\frac{5}{2}-\frac{3}{2 x} \cdot z \\
\Rightarrow \frac{d z}{d x}+\frac{3}{2 x} \cdot z=-\frac{5}{2}"

"I \cdot F=e^{\int \frac{3}{2 x} d x}=e^{\frac{3}{2} \cdot \ln x}=e^{\ln x^{3 / 2}}=x^{3 / 2}"

Multiplying above differential equation by I . F, and integrating,

"\Rightarrow z \cdot x^{3 / 2}=\int -\frac{5}{2}.x^\frac{3}{2}dx+c\\ 
\Rightarrow z \cdot x^{3 / 2}=-\frac{5}{2} \cdot \frac{x^{\frac{3}{2}+1}}{\frac{3}{2}+1}+c \\
\Rightarrow z \cdot x^{3 / 2}=-\frac{5}{2} \cdot \frac{2}{5} \cdot x^{\frac{5}{2}}+c\\
\Rightarrow e^{y} \cdot x^{3 / 2}=-x^{\frac{5}{2}}+c\left[\because z=e^{y}\right]\\"

"\Rightarrow e^{y}=-x^{\frac{5}{2}-\frac{3}{2}}+c. x^{-3 / 2}\\
\Rightarrow e^{y}=-x+c x^{-\frac{3}{2}}\\
\Rightarrow e^{y}=c x^{-\frac{3}{2}}-x\\
\Rightarrow y=ln|c x^{-\frac{3}{2}}-x|"